OpenStudy (anonymous):

The derivative of 8^x+8*7^x why do u get ln8+8x...

OpenStudy (anonymous):

8^x*

OpenStudy (amistre64):

most likely it is a consequence of way it is worked out ....

OpenStudy (amistre64):

yousually there is a ln(y=b^x) thing going on to ease the process

OpenStudy (anonymous):

So anything raised to the x would be like this 6*ln(6)

OpenStudy (amistre64):

almost, but sure

OpenStudy (experimentx):

y = a^x where 'a' is be any constant dy/dx = ln(a) a^x

OpenStudy (experimentx):

best done using implicit differentiation let y = a^x taking log both sides. ln(y) = x ln(a) 1/y dy/dx = ln(a) => dy/dx = y ln(a) => a^x ln(a)

OpenStudy (experimentx):

the other way is to change a to e let a = e^u ln(a) = u ln(e) = u therefore we have y = e^(ln(a)x) => use chain rule here, you will get the same result.

OpenStudy (experimentx):

*** a = e^ln(a)